OpenStudy (anonymous):

Solve using substitution: 2y − x + 2 = 0 -3x + 8y = -5 a. (-10, 4) b. (3, 1/2) c. (2, 0) d. (4, 1)

OpenStudy (anonymous):

b. (3, 1/2)

OpenStudy (lgbasallote):

2y - x +2 = 0 add x toboth sides 2y + 2 = x now substitute this into the x of the second equation -3(2y +2) + 8y = -5 -6y - 6 + 8y = -5 2y - 6 = -5 2y = -5 + 6 2y = 1 do you get it?